OpenStudy (rainbowdashrocks):

Joseph wants to estimate 1 3/7 * 5 4/5 . To estimate, he simplifies the expression 1 1/2 *6 Is Joseph’s estimate an underestimate or an overestimate? A.underestimate B.overestimate

OpenStudy (rainbowdashrocks):

will fan and medal plz help

OpenStudy (anonymous):

1 1/2 *6=9 and =\[8\frac{2 }{ 7 }\] So what do you think?

OpenStudy (anonymous):

1 3/7 * 5 4/5 is the one that = 8 2/7

OpenStudy (anonymous):

Would you think its a over estimate or a under?

OpenStudy (rainbowdashrocks):

b?

OpenStudy (anonymous):

I believe that is correct! :)

OpenStudy (rainbowdashrocks):

Thx